The Moon company currently sells for 100 $. The annual stock price volatility is %10 and risk- free interest rate %8, the price of a call on a company’s stock with strike price 200 $ and time period 2 months. Find the stock option price with Black and Scholes Model. If option market value 320 $ what is the option strategy?


1
Expert's answer
2021-02-03T01:23:46-0500

"c=S0\times N(d1)-K\times N(d2)\times DF""d1=\frac{ln(\frac{S0}{K})+(r+\frac{\sigma^2}{2})\times t}{ \sigma\times t^{0.5}}"


"d2=d1-\sigma\times t^{0.5}"


"DF=e^{-r\times t}"

S0=100

K=200

"\sigma=0.1"

r=0.08, 0.0067 years

t=2, 0.167 years

"DF=e^{-0.0067\times0.167}=0.999"

"d1=\frac{ln(\frac{100}{200})+(0.0067+\frac{0.1^2}{2})\times 0.167}{ 0.1\times0.167^{0.5}}=-0.6455"

"d2=-0.6455-0.1\times 0.167^{0.5}=0.6863"

"c=100\times N(-0.6455)-200\times N(-0.6863)\times0.999=-23.27"


K=320

"DF=e^{-0.0067\times0.167}=0.999"

"d1=\frac{ln(\frac{100}{320})+(0.0067+\frac{0.1^2}{2})\times 0.167}{ 0.1\times0.167^{0.5}}=-1.1155"

"d2=-1.1155-0.1\times 0.167^{0.5}=-1.1563"

"c=100\times N(-1.1155)-320\times N(-1.1563)\times0.999=-26.33"